# Analyticity of the Black Hole S-Matrix

### v1

@@ Appendix A Analyticity for compact and exponentially decaying potentials
There is also analyticity in the lower half-plane. For $\mathrm{Im}\,\omega<0$ the integrand of (30) instead may now diverge for $x^{\prime}\to$-$\infty$ where it behaves as $\sim e^{(2\,|\mathrm{Im}\,\omega|-\mu_{-})\,|x^{\prime}|}$ (recall that $\chi_{L}(\omega,x\to-\infty)\to 1$ according to the boundary condition (15) where $\chi_{L}(\omega,x)=e^{i\omega x}\phi_{L}(\omega,x)$). This implies that $\mathrm{Im}\,\omega<\mu_{-}/2$ in order to have convergence. [^27] Note that the opposite limit $x^{\prime}\to+\infty$ is safe: the exponential growth of $\chi_{L}(\omega,x)$ from (130) is canceled by the pre-factor $e^{-2i\omega x^{\prime}}\sim e^{-2|\mathrm{Im}\,\omega|x^{\prime}}$ in the integrand. We thus establish that
$A_{\text{out}}(\omega)\;\text{ is analytic for }\;{\mu_{-}\over 2}>\mathrm{Im}\,\omega>-{\mu_{+}\over 2}.$ (138)
@@
$R(\omega)\;\text{ is meromorphic for }\;{\mu_{-}\over 2}>\mathrm{Im}\,\omega>-{\min(\mu_{-},\mu_{+})\over 2}.$ (140)
@@ Appendix D Bootstrap bounds on Love numbers without gravity
We assume that the potential is such that the reflection coefficient $R(\omega)$ is analytic in the UHP. Additionally, unitarity implies that $|R(\omega)|\leq 1$ for $\omega\in\mathbb{R}$. Hence, by the maximum modulus principle, $|R(\omega)|\leq 1$ for all values of $\omega$ in the UHP. later
@@
By assuming the reflection coefficient is analytic and bounded, we have constrained the first two Love numbers in our worldline EFT. Higher-order coefficients could be computed using the procedure set out in EliasMiro:2019kyf; EliasMiro:2021nul.

### v2

@@ Appendix A Analyticity for compact and exponentially decaying potentials
There is also analyticity in the lower half-plane. For $\mathrm{Im}\,\omega<0$ the integrand of (30) instead may now diverge for $x^{\prime}\to$-$\infty$ where it behaves as $\sim e^{(2\,|\mathrm{Im}\,\omega|-\mu_{-})\,|x^{\prime}|}$ (recall that $\chi_{L}(\omega,x\to-\infty)\to 1$ according to the boundary condition (15) where $\chi_{L}(\omega,x)=e^{i\omega x}\phi_{L}(\omega,x)$). This implies that $|\mathrm{Im}\,\omega|<\mu_{-}/2$, or $\mathrm{Im}\,\omega>-\mu_{-}/2$, in order to have convergence. [^30] Note that the opposite limit $x^{\prime}\to+\infty$ is safe: the exponential growth of $\chi_{L}(\omega,x)$ from (130) is canceled by the pre-factor $e^{-2i\omega x^{\prime}}\sim e^{-2|\mathrm{Im}\,\omega|x^{\prime}}$ in the integrand. We thus establish that
$A_{\text{out}}(\omega)\;\text{ is analytic for }\;{\mu_{+}\over 2}>\mathrm{Im}\,\omega>-{\mu_{-}\over 2}.$ (138)
@@
$R(\omega)\;\text{ is meromorphic for }\;{\mu_{+}\over 2}>\mathrm{Im}\,\omega>-{\min(\mu_{-},\mu_{+})\over 2}.$ (140)
@@ Appendix D Bootstrap bounds on Love numbers without gravity
We assume that the potential is such that the reflection coefficient $R(\omega)$ is analytic and that $|R(\omega)|$ is bounded in the UHP. Additionally, unitarity implies that $|R(\omega)|\leq 1$ for $\omega\in\mathbb{R}$. Hence, by the Phragmén-Lindelöf principle ahlfors1973, $|R(\omega)|\leq 1$ for all values of $\omega$ in the UHP.
@@
By assuming the reflection coefficient is analytic and bounded, we have constrained the first two Love numbers in our worldline EFT. Higher-order coefficients could be bounded using the procedure set out in EliasMiro:2019kyf; EliasMiro:2021nul.